OpenStudy (anonymous):

When the shuttle bus comes to a sudden stop to avoid hitting a dog, it decelerates uniformly at 5.0 m/s 2 as it slows from 9.5 m/s to 0 m/s. Find the time interval of acceleration for the bus. Answer in units of s

OpenStudy (anonymous):

a = Δ v/ Δt -> Δt = Δv/ a = -9.5 /- 5 = ...

OpenStudy (anonymous):

given acceleration =-5m/s^2 velocity slow down from v2=9.5 m/s to v1= 0 m/s if the formula for distance is s= change in V/ change in t change in t=change in V/s t= (0-9.5)/-5 t= ?